# Enhancing Vision-Language Models with Scene Graphs for Traffic Accident Understanding

### v1

@@ Abstract
Recognizing a traffic accident is an essential part of any autonomous driving or road monitoring system. An accident can appear in a wide variety of forms, and understanding what type of accident is taking place may be useful to prevent it from reoccurring. The task of being able to classify a traffic scene as a specific type of accident is the focus of this work. We approach the problem by likening a traffic scene to a graph, where objects such as cars can be represented as nodes, and relative distances and directions between them as edges. This representation of an accident can be referred to as a scene graph, and is used as input for an accident classifier. Better results can be obtained with a classifier that fuses the scene graph input with representations from vision and language. This work introduces a multi-stage, multimodal pipeline to pre-process videos of traffic accidents, encode them as scene graphs, and align this representation with vision and language modalities for accident classification. When trained on 4 classes, our method achieves a balanced accuracy score of 57.77% on an (unbalanced) subset of the popular Detection of Traffic Anomaly (DoTA) benchmark, representing an increase of close to 5 percentage points from the case where scene graph information is not taken into account.
@@ I Introduction
The task of understanding traffic scenarios is one with ever-increasing importance, particularly for the advancement of Autonomous Vehicle (AV) and road infrastructure systems (Xu et al., 2021; Yao et al., 2022; Qasemi et al., 2023; Francis et al., 2022). A major aspect of this task is to efficiently and accurately recognize different types of traffic accidents, with the ultimate goal of preventing them. In this work, we approach this challenge by modeling traffic scenes using scene graphs. To improve classification performance, the graph representation is further aligned with representations from the vision and language modalities within contrastively-trained foundation models. Starting with video clips of traffic incidents, procured from the Detection of Traffic Anomaly (DoTA) dataset Yao et al. (2022), we build Scene-Traffic-Graph Inference (STGi), a multi-stage, multimodal pipeline to pre-process videos, encode them as scene graphs, and align these representations with vision and language modalities to classify traffic accidents.
In this paper, we propose Scene-Traffic-Graph Inference(STGi) as a unified system for accident classification. First, we generate scene graphs using the roadscene2vec (rs2v) tool Malawade et al. (2021). Then, we use a scene graph encoder (SGE) from the same authors Malawade et al. (2022) to obtain encodings that are aligned to encodings of textual and visual inputs, obtained from CLIP Radford et al. (2021) and X-CLIP Ni et al. (2022), respectively. X-CLIP directly expands upon CLIP’s image encoder to include an attention mechanism to model inter-frame communication and to generate a new embedding representation from video frames. In essence, scene graphs are treated as a new modality (or ‘view’) that is aligned with text and video signals, and are all used together to classify traffic accident scenes.
A summary of contributions is as follows:
  • We introduce Scene-Traffic-Graph Inference (STGi), a novel method for traffic accident classification which leverages scene graphs to capture the essential features of a traffic accident.
  • We demonstrate that the added signal from a scene graph modality can enhance the performance of a video-language traffic accident classifier by nearly 5 percentage points.
  • Experiments in this work demonstrate that aligning the scene graph modality with vision and language give similar results to omitting alignment training, although increasing the batch size and training time during alignment shows a trend of increasing scores and the potential in further improving classification results.
@@ II-A Scene Graphs for Representing Traffic Scenes
Several works have taken the approach of modeling traffic scenarios using scene graphs (SGs). Yu et al. (2020) propose one such representation that forms the basis of the SGs produced in our work. They define a graph structure with nodes representing entities such as vehicles and pedestrians while edges represent relations between the objects and are labeled by distance category (e.g., visible, very near) and orientation (e.g., in front of, beside). Zipfl and Zöllner (2021) also model these objects in a traffic scene, further accounting for relative speeds of objects in the graph edges. Guo et al. (2023) extracts similar semantic information from a traffic scene, but the focus is on generating a knowledge graph, accounting for added traffic information present in a scene such as road signs and lanes. Qasemi et al. (2023) also use a knowledge graph to enhance a video-language model for the related task of Question Answering in the domain of traffic accidents. Their method is to inject domain-specific knowledge by first auto-generating video annotations and storing semantic information about a traffic scene in a knowledge graph. The use of a knowledge graph in this case is helpful in contextualizing the traffic scene in a broader traffic monitoring ontology and for answering general traffic accident-related questions, such as the impact of having more vehicles in a traffic scene, but is less effective in identifying accident types based on the interactions among vehicles in a specific traffic scene. Instead, our method is better suited for the task of accident classification by leveraging scene graphs to capture relationships among objects that may be involved in a traffic accident in a given video example. Combining the use of scene graphs with knowledge graphs to model traffic settings, Dimasi (2023) makes use of a knowledge graph to boost the performance of the RelTR (Relation Transformer) model Cong et al. (2022), an encoder-decoder model used for scene graph generation by framing it as a set prediction problem.
@@ II-B Leveraging Scene Graphs with Neural Networks
There are many possible designs and applications for leveraging the versatile structure of scene graphs to enhance the performance of neural networks Li et al. (2019); Khademi and Schulte (2020); Rana et al. (2023). More generally, Zhang et al. (2019) describe the ability of scene graphs to enhance Visual Question Answering (VQA) systems as an alternative to solely focusing on vision and language features. Although their study is not specific to our traffic domain, they claim that scene graphs derived from images can capture the essential visual features and may outperform images for VQA tasks. Many works have focused on automatically generating scene graphs that can easily be encoded into neural networks for various downstream tasks. The rs2v tool Malawade et al. (2021) generates scene graphs for individual frames of traffic videos and is applied to our work. Alternatively, Nag et al. (2023) design a tool specifically for modeling videos with scene graphs which accounts for spatial and temporal relationships between frames based on sequence modeling with transformers. However, they train on the Action Genome Ji et al. (2019) dataset and produce graphs that are unrelated to traffic scenarios, making this tool less applicable to our work.
@@ II-C Multimodal Alignment with Contrastive Representations
Radford et al. (2021) proposed a model that uses a contrastive training strategy to align text captions with images for the Contrastive Language–Image Pre-training (CLIP) model. From this, several works have focused on aligning encoders of additional modalities to those trained for CLIP, such as aligning audio Wu et al. (2022) and haptic data Tatiya et al. (2024). Similar to our implementation, Koch et al. (2023) use CLIP’s text encoder to contrastively train 3D scene graphs. The work by Huang et al. (2023) aims to enhance CLIP’s performance by injecting scene graph information into the embedding generated by CLIP’s text encoder before contrastive learning is performed between the image and text embedding spaces. To the best of our knowledge, however, our work is the first to treat scene graphs as an additional modality that is grounded to a text and video encoder. There are also many possible strategies for fusing the outputs of a multimodal model before the downstream task at hand, such as early and late stage fusion involving concatenation Sleeman et al. (2022); Pawłowski et al. (2023), merging Sleeman et al. (2022), or sampling from a shared embedding space Gao et al. (2020). We experiment with several of these late fusion techniques on the three modality embeddings after training for alignment.
@@ III Methodology
To approach the problem of traffic accident classification, a multi-stage pipeline was developed. The pipeline, which is represented in Figure 1, can be summarized in four stages and is described in the following sections:
@@ III-A Data Pre-processing
In the first stage, for extensibility and ease, we leverage and tune existing tools for generating traffic scene graphs that will later be fed into our modeling approach. Various scene graph generators (SGGs) are available for this, although not many fit our specific requirements for a generator that is capable of identifying and encoding features unique to traffic accidents, such as different types of vehicle collisions. Some do work for our needs — in particular, the roadscene2vec (rs2v) tool created by Malawade et al. (2021) which we leverage for generating scene graphs from traffic video frames. To use rs2v, we first sample a fixed amount of frames from each video, and text captions are generated manually for each class. Then, feeding in a series of frames from a traffic scene, the rs2v generator uses an object detector and keeps only the relevant entities relating to traffic (such as road, cars, pedestrians, etc.), filtering out other detected objects. Next, after generating a “bird’s eye view” (BEV) projection of the image and approximating the relative location of each object in this projection, edges are connected between nearby entities in the scene, and vehicles are mapped to lanes. This forms the scene graph for a specific frame from a video sample. The SGG is an important component of this pipeline as it defines the elements in the scene graph modality.
@@
An important step before employing this tool is to calibrate the BEV and adjust the proximity thresholds (which are used to create edges of varying attributes, such as very near or visible, relating a pair of objects). The authors specifically note the challenges they faced when adjusting these settings with regard to the DoTA dataset, and mention that for best results, they needed to calibrate their model for each traffic scene Malawade et al. (2022). In our case, we iteratively sample and adjust the BEV parameters and proximity thresholds based on the output quality of the SGs produced for various scenes in DoTA. We do this to select one configuration to generate the SGs for our task, although it remains an inherent challenge to generalize the parameter settings for the entire dataset. Some examples of the generated SGs can be seen in Appendix A.
@@
For text, we manually compose captions describing each of the four accident classes (see Appendix B), pairing each caption with videos from its respective class to form the training examples. We also experiment with two sets of captions, however, we focus the scope of this work primarily on the scene graph generation process from video frames rather than caption generation. It should be noted that although these captions are used during training for aligning the SGE, they are not used during inference on the final model because they are not from our dataset and provide a one-to-one mapping directly to the accident classes. Through fine-tuning the classification head using these captions, the model can achieve 100% accuracy this way. Instead, we fine-tune and test the classification head using the same caption as noted in Appendix B for all examples.
@@ III-B Scene Graph Encoder
In stage two, the SGE encodes scene graphs to fixed-length embeddings. To do this, a multirelational graph convolutional network (MRGCN) as employed by Malawade et al. (2022) is used, which includes an attention mechanism along with LSTMs to model the spatial and temporal relations of the scene graphs generated for a given video. The SGE may be pre-trained for the classification task before aligning the scene graphs with the language and vision modalities.
@@ IV-A Dataset
Given the nature of our multimodal approach, we require data in the form of videos, text, and scene graphs. Regarding the videos, we focus on using the Detection of Traffic Anomaly (DoTA) dataset Yao et al. (2022) for this task. DoTA is a dataset consisting of 4,677 curated videos of traffic anomalies from YouTube, and was specifically designed to answer the following three questions about each video: (1) When does the anomalous event (i.e., accident) start and end; (2) Where are the anomalous regions of each video frame; and (3) What type of anomaly is taking place. For our use case of this dataset, question one is irrelevant since we only sample frames from the anomalous portion of the video based on the annotations of the DoTA authors. Additionally, question two is less essential for us since we use a SGG, which is meant to extract the necessary objects from video frames and model their relevant spatial relationships for our task. We therefore prioritize providing an answer for the third question, that is, classifying different traffic anomalies. However, given the multitude of classes and the inherent challenge in distinguishing them even with the naked eye, we limit the problem to only a subset of four of the available classes.
In addition to the provided video classes, Yao et al. (2022) also divides the videos into two types, ego (where the vehicle from which the video recording is captured is involved in the accident), and its opposite, non-ego. Every video class contains videos of both types. By default, rs2v is designed to only model relationships between the ego vehicle and other objects, not among other objects themselves, so we further restrict our task to the set of ego videos available in DoTA. In total, this leaves us with 4 classes and 2,163 videos. The classes are labeled as follows, describing the type movement of the vehicle as the accident occurs: moving_ahead_or_waiting, oncoming, turning, and lateral.
@@ IV-B Baselines and Metrics
In order to evaluate the usefulness of representing traffic scenes graphically to classify them, we first verify whether a SGE alone is superior to randomly guessing an accident class. We later evaluate the effectiveness of using SGs in a multimodal classifier by comparing the difference in the performance between using just text and video embeddings versus adding in the signal from the SG modality.
For this and other results, the primary metrics we focus on are accuracy and balanced accuracy. Accuracy is simply defined as the proportion of the correctly classified predictions out of the total number of predictions and remains an interpretable metric that is relatively simple to compare across different benchmarks. There is, however, significant imbalance in the dataset, with the turning class appearing over twice as frequently as the remaining three. Since we want to evenly prioritize classifying each type of accident, we turn to balanced accuracy, which is computed as the (unweighted) average recall over all classes, as a more reliable metric for our use case.
IV-C Modality Alignment, Pre-training and Hyperparameters
The core TG model consists of three encoders for three different modalities: text, vision, and graph. For our classification task, a MLP classifier head is placed at the end of the model to output scores for each of the four classes. The text and vision encoders are both kept frozen throughout all of our experiments as we train the SGE to align its embedding space with those of text and vision using Symmetric Cross Entropy Loss Wang et al. (2019). To evaluate the effectiveness of this training regiment, we analyze the classification performance with and without training for modality alignment. We also experiment with the effect of pre-training the SGE before conducting the alignment training. As an additional pre-training baseline, we use a pre-trained model provided by Malawade et al. (2022), which was trained on a synthetic traffic dataset generated using the CARLA traffic simulator Dosovitskiy et al. (2017). Finally, we perform tests to evaluate the performance of different hyperparameter settings (batch size, training time) for the alignment training.

### v2

@@ Abstract
Recognizing a traffic accident is an essential part of any autonomous driving or road monitoring system. An accident can appear in a wide variety of forms, and understanding what type of accident is taking place may be useful to prevent it from reoccurring. This work focuses on classification of traffic scenes into specific accident types. We approach the problem by representing a traffic scene as a graph, where objects such as cars can be represented as nodes, and relative distances and directions between them as edges. This representation of a traffic scene is referred to as a scene graph, and can be used as input for an accident classifier. Better results are obtained with a classifier that fuses the scene graph input with visual and textual representations. This work introduces a multi-stage, multimodal pipeline that pre-processes videos of traffic accidents, encodes them as scene graphs, and aligns this representation with vision and language modalities before executing the classification task. When trained on 4 classes, our method achieves a balanced accuracy score of 57.77% on an (unbalanced) subset of the popular Detection of Traffic Anomaly (DoTA) benchmark, representing an increase of close to 5 percentage points from the case where scene graph information is not taken into account.
@@ I Introduction
The task of understanding traffic scenarios is one with ever-increasing importance, particularly for the advancement of Autonomous Vehicles (AV) and road infrastructure systems (Xu et al., 2021; Yao et al., 2022; Qasemi et al., 2023; Francis et al., 2022; Huang et al., 2024; Stoler et al., 2024a, b). A major aspect of this task is to efficiently and accurately recognize different types of traffic accidents, with the ultimate goal of preventing them. We approach this challenge by modeling traffic scenes using scene graphs. To improve classification performance, the graph representation is further aligned with representations from the vision and language modalities within contrastively-trained foundation models. Starting with video clips of traffic incidents, procured from the Detection of Traffic Anomaly (DoTA) dataset Yao et al. (2022), we build Scene-Traffic-Graph Inference (STGi), a multi-stage, multimodal pipeline that pre-processes videos, encodes them as scene graphs, and aligns these representations with vision and language modalities before classifying traffic accidents.
We propose STGi as a unified system for accident classification. To start, in STGi we generate scene graphs using the roadscene2vec (rs2v) tool Malawade et al. (2021). Then, we use a scene graph encoder (SGE) from the same authors Malawade et al. (2022) of rs2v to obtain encodings that are aligned to encodings of textual and visual inputs, obtained from CLIP Radford et al. (2021) and X-CLIP Ni et al. (2022), respectively [^1] CLIP is a multi-modal model that is capable of zero-shot image classification by measuring text-image similarity, and its main components are a text encoder and an image encoder. X-CLIP expands upon CLIP’s image encoder to include an attention mechanism to model inter-frame communication and to generate a new embedding representation from video frames.. Scene graphs are treated as a new modality (or ‘view’) that is aligned with text and video signals, and all are used together to classify traffic accident scenes. A summary of contributions is as follows:
  • We introduce Scene-Traffic-Graph Inference (STGi), a novel classification method that leverages scene graphs to capture essential features from traffic accidents.
  • We show that the added signal from a scene graph modality can enhance the performance of a video-language traffic accident classifier by nearly 5 percentage points.
  • Experiments in this work demonstrate that aligning the scene graph modality with vision and language gives similar results to omitting alignment training, although increasing the batch size and training time during alignment shows a trend of increasing scores and the potential in further improving classification results.
@@ II-A Scene Graphs for Representing Traffic Scenes
Several works have taken the approach of modeling traffic scenarios using scene graphs (SGs). Yu et al. (2020) propose one such representation that forms the basis of the SGs produced in our work. They define a graph structure with nodes representing entities such as vehicles and pedestrians while edges represent relations between the objects and are labeled by distance category (e.g., visible, very near) and orientation (e.g., in front of, beside). Zipfl and Zöllner (2021) also model these objects in a traffic scene, further accounting for relative speeds of objects in the graph edges. Guo et al. (2023) extract semantic information from a traffic scene as well, but their focus is largely on generating graphs from road signs and lanes to understand how a traffic scene may be interpreted from the perspective of a driver. Qasemi et al. (2023) use a knowledge graph (KG) to enhance a video-language model for the related task of Question Answering in the domain of traffic accidents. Their method is to inject domain-specific knowledge by first auto-generating video annotations and storing semantic information about a traffic scene in a KG. It is interesting to contrast their KG-based approach, which helps contextualize the traffic scene in a broader traffic monitoring ontology, with our scene graph method, which builds a unique graphical structure from each scene. In the former case, the KG can answer general traffic accident-related questions, such as the impact of having more vehicles in a traffic scene, but is less effective in identifying accident types based on the interactions among vehicles in a specific traffic scene. Instead, our method is better suited for the task of accident classification by leveraging SGs to capture relationships among objects that may be involved in a traffic accident in a given video example. Combining the use of SGs with KGs to model traffic settings, Dimasi (2023) makes use of a KG to boost the performance of the RelTR (Relation Transformer) model Cong et al. (2022), an encoder-decoder model used for scene graph generation by framing it as a set prediction problem.
@@ II-B Leveraging Scene Graphs with Neural Networks
There are many possible designs and applications for leveraging the versatile structure of SGs to enhance the performance of neural networks Li et al. (2019); Khademi and Schulte (2020); Rana et al. (2023); Saxena et al. (2024). Zhang et al. (2019) describe the ability of SGs to enhance Visual Question Answering (VQA) systems as an alternative to solely focusing on vision and language features. Although their study is not specific to our traffic domain, they claim that SGs derived from images can capture the essential visual features and may outperform images for VQA tasks. Many works have focused on automatically generating SGs that can easily be encoded into neural networks for various downstream tasks. For example, the rs2v tool Malawade et al. (2021) generates SGs for individual frames of traffic videos and is applied to our work. Alternatively, Nag et al. (2023) design a tool specifically for modeling videos with SGs which accounts for spatial and temporal relationships between frames based on sequence modeling with transformers. However, they train on the Action Genome dataset Ji et al. (2019) and produce graphs that are unrelated to traffic scenarios, making this tool less applicable to our work.
@@ II-C Multimodal Alignment with Contrastive Representations
Radford et al. (2021) propose a model that uses a contrastive training strategy to align text captions with images for the Contrastive Language–Image Pre-training (CLIP) model. From this, several works have focused on aligning encoders of additional modalities to those trained for CLIP, such as audio Wu et al. (2022) and haptic modalities Tatiya et al. (2024). Similar to our implementation, Koch et al. (2023) use CLIP’s text encoder to contrastively train 3D scene graphs, which are graphs that implicitly contain spatial information of objects in a scene, and can also model other relationships between these objects Koch et al. (2023). The work by Huang et al. (2023) aims to enhance CLIP’s performance by injecting scene graph information into the embedding generated by CLIP’s text encoder before contrastive learning is performed between the image and text embedding spaces. To the best of our knowledge, however, our work is the first to treat scene graphs as an additional modality that is grounded to a text and video encoder. There are also many possible strategies for fusing the outputs of a multimodal model before the downstream task at hand, such as early and late stage fusion involving concatenation Sleeman et al. (2022); Pawłowski et al. (2023), merging Sleeman et al. (2022), or sampling from a shared embedding space Gao et al. (2020). We experiment with several of these late fusion techniques on the three modality embeddings after training for alignment.
@@ III Methodology
To approach classification of traffic accidents, a multi-stage pipeline was developed. The pipeline, which is represented in Figure 1, can be summarized in four stages that are described in the following sections:
@@ III-A Data Pre-processing
In the first stage, for extensibility and ease, we leverage and tune existing tools for generating traffic scene graphs that will later be fed into our modeling approach. Various scene graph generators (SGGs) are available for this, although not many fit our specific requirements for a generator that is capable of identifying and encoding features unique to traffic accidents, such as different types of vehicle collisions. Some do work for our needs — in particular, the roadscene2vec (rs2v) tool created by Malawade et al. (2021) that we leverage for generating scene graphs from traffic video frames. To use rs2v, we first sample a fixed amount of frames from each video, and for each class we manually generate text captions. Then, feeding in a series of frames from a traffic scene, the rs2v generator uses an object detector and keeps only the relevant entities relating to traffic (such as road, cars, pedestrians, etc.), filtering out other detected objects. Next, after generating a “bird’s eye view” (BEV) projection of the image and approximating the relative location of each object in this projection, edges are added to connect nearby entities in the scene and to map vehicles to lanes. This forms the scene graph for a specific frame from a video sample. The SGG is a crucial component of this pipeline as it defines the elements in the scene graph modality.
@@
An important step before employing this tool is to calibrate the BEV and adjust the proximity thresholds (which are used when creating edges labeled with varying attributes, such as very near or visible, relating a pair of objects). The authors specifically note the challenges they faced when adjusting these settings with regard to the DoTA dataset, and mention that for best results, they needed to calibrate their model for each traffic scene Malawade et al. (2022). In our case, we generate SGs for a random subset of DoTA traffic scenes, visually inspect them for errors (e.g., the SGG labels the edge relating the ego vehicle to a vehicle directly in front of it as visible rather than very near), and then manually adjust the BEV parameters and proximity thresholds in an iterative manner. We do this to select one configuration to generate the SGs for our task, although it remains an inherent challenge to generalize the parameter settings for the entire dataset. Some examples of the generated SGs can be seen in Appendix A.
@@
For text, we manually compose captions describing each of the four accident classes (see Appendix B), and pair each caption with videos from its respective class to form the training examples. We also experiment with two sets of captions, however, we focus the scope of this work primarily on the scene graph generation process from video frames rather than caption generation. It should be noted that although these captions are used during training for aligning the SGE, they are not used during inference on the final model because they are not from our dataset and provide a one-to-one mapping directly to the accident classes. Through fine-tuning the classification head using these captions, the model can achieve 100% accuracy. Instead, we fine-tune and test the classification head using the same caption, as noted in Appendix B, for all examples.
@@ III-B Scene Graph Encoder
In stage two, the SGE encodes sequences of scene graphs derived from a given video to fixed-length embeddings. To do this, a multirelational graph convolutional network (MRGCN) as employed by Malawade et al. (2022) is used, which includes an attention mechanism along with LSTMs to model the spatial and temporal relations of the scene graphs generated for a given video. The SGE may be pre-trained for the classification task before aligning the scene graphs with the language and vision modalities.
@@ IV-A Dataset
Our multimodal approach requires data in the form of videos, text, and scene graphs. We focus on using the Detection of Traffic Anomaly (DoTA) dataset Yao et al. (2022) for videos for this task. DoTA is a dataset of 4,677 curated videos of traffic anomalies from YouTube, and was specifically designed to answer the following three questions about each video: (1) When does the anomalous event (i.e., accident) start and end; (2) Where are the anomalous regions of each video frame; and (3) What type of anomaly is taking place. For our use case, question one is irrelevant since we only sample frames from the anomalous portion of the video based on the annotations of the DoTA authors. Question two is also less essential for us since we use a SGG, which extracts the necessary objects from video frames and models their relevant spatial relationships. We therefore prioritize answering the third question, that is, classifying different traffic anomalies. However, given the multitude of classes and the inherent challenge in distinguishing them even with the naked eye, we limit the problem to a subset of four of the available classes.
In addition to the video classes, Yao et al. (2022) also divide the videos into two types (which are different from classes), ego (where the vehicle from which the video recording is captured is involved in the accident), and its opposite, non-ego. Every video class has videos of both types (e.g., there are both ego and non-ego videos that fall into the turning class). By default, rs2v is designed to only model relationships between the ego vehicle and other objects, not among other objects themselves, so we further restrict our task to the set of ego videos available in DoTA. In total, this leaves us with four classes and 2,163 videos. The classes labels, describing the vehicle movement as the accident occurs, are as follows: moving_ahead_or_waiting, oncoming, turning, and lateral.
@@ IV-B Baselines and Metrics
In order to evaluate the usefulness of representing traffic scenes graphically for classification, we verify whether a SGE alone is superior to randomly guessing an accident class. We then assess the effectiveness of SGs in a multimodal classifier by comparing the difference in performance between using just text and video embeddings versus including the SG modality. Regarding metrics, we focus on accuracy and balanced accuracy. Accuracy is simply defined as the proportion of the correctly classified predictions out of the total number of predictions and is an interpretable metric that can be easily compared across benchmarks. However, there is significant imbalance in the dataset, with the turning class appearing over twice as frequently as the other three. We use balanced accuracy, computed as the (unweighted) average recall over all classes, to give equal priority to each accident class.
IV-C Modality Alignment, Pre-training, and Hyperparameters
The core STGi model consists of three encoders for three different modalities: text, vision, and graph. For our classification task, a MLP classifier head is placed at the end of the model to output scores for each of the four classes. The text and vision encoders are both kept frozen throughout all of our experiments as we train the SGE to align its embedding space with those of text and vision using Symmetric Cross Entropy Loss Wang et al. (2019). To evaluate the effectiveness of this training regiment, we analyze the classification performance with and without training for modality alignment. We also experiment with the effect of pre-training the SGE before conducting the alignment training. As an additional pre-training baseline, we use a model provided by Malawade et al. (2022), which was pre-trained on a synthetic traffic dataset generated using the CARLA traffic simulator Dosovitskiy et al. (2017). Finally, we conduct tests to evaluate the performance of different hyperparameter settings (batch size, training time) for the alignment training.